OpenStudy (anonymous):

How come arctan(-sqrt(3)) doesn't = -pi/6?

OpenStudy (anonymous):

If it equals -pi/3 then\[\frac{ \frac{ -1 }{ 2 } }{ \frac{ \sqrt(3) }{ 2 } }~=~\frac{ -1 }{ \sqrt(3) }*\frac{ \sqrt(3) }{ \sqrt(3) }~=\frac{ -\sqrt(3) }{ 3 }\]

OpenStudy (anonymous):

My calculator tells me it is -pi/3.

myininaya (myininaya):

it should be -pi/3 since at -pi/3 cosine has value 1/2 and sine value -sqrt(3)/2 tangent=sine/cosine

myininaya (myininaya):

\[\tan(- \frac{\pi}{3})=\frac{\sin(-\frac{\pi}{3})}{\cos(-\frac{\pi}{3})}=\frac{\frac{-\sqrt{3}}{2}}{\frac{1}{2}}=\frac{-\sqrt{3}}{1}=-\sqrt{3} \\ \arctan(-\sqrt{3})=\frac{- \pi }{3}\]

OpenStudy (anonymous):

http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_sm.gif -pi/3 is ( sqrt(3)/2, -1/2 )?

OpenStudy (anonymous):

cosine is x sine is y

myininaya (myininaya):

you are looking at the wrong point

myininaya (myininaya):

it isn't (sqrt(3)/2,-1/2) at -pi/3

OpenStudy (anonymous):

Am I looking at -pi/6?

myininaya (myininaya):